Solve the equation if a²+b²+c²=2(a-b-c)-3 then find out the value of 2a-3b+4c=?
 

Grade:12th pass

2 Answers

jagdish singh singh
7 years ago
\hspace{-0.7 cm}$ we can write $a^2+b^2+c^2=2(a-b-c)-3$\\\\ as $a^2-2a+1+b^2+2b+1+c^2+2c+1=0$\\\\ So $(a-1)^2+(b+1)^2+(c-1)^2=0$\\\\ Using sum of square of all three real quantity is zero, if all are zero\\\\ So $a-1 = 0$ and $b+1 = 0$ and $c+1=0$\\\\ So $a=1\;\;,b=-1\;\;,c=-1$. So $2a-3b+4c =2+3-4=1 $
ankit singh
3 years ago
a² + b² + c² = 2 (a - b - c ) - 3
=> a² + b² + c² = 2a - 2b - 2c -3
=> a² + b² + c² - 2a + 2b + 2c + 3 = 0
//rearrange the above as below
=> a² - 2a + 1 + b² + 2b + 1  + c² + 2c + 1 = 0
=> (a-1)² + (b+1)² + (c+1)² = 0
Using sum of all three real quantity is zero, if all are zero
=>  a - 1  = 0 => a = 1
     b + 1  = 0  => b = -1
     c + 1 = 0 => c = - 1
Thus value of 2a - 3b + 4c = 2 * 1 - 3(-1) + 4(-1) = 1